OpenStudy (anonymous):

Your uncle is making you a desk for you to take to college. One side of the desk is 32 inches and the diagonal length is 36 inches. What is the length of the other side of the desk?

OpenStudy (anonymous):

is 32 the longest lenght?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

a^2 + b^2 = c^2 a^2 = c^2 - b^2

OpenStudy (anonymous):

The answer is 35.

OpenStudy (anonymous):

yea i got that too

OpenStudy (anonymous):

its either 13.5 inches 27.2 inches 17.4 in 16.5 inches.